OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x))

OpenStudy (zzr0ck3r):

\[f(x) = 9x-2\\g(x) = -x+3=3-x\\f(g(x)) = f(3-x) = 9(3-x)-2 = ?\]just simplify from there